### v2

@@ 1 Introduction
A central problem in scientific machine learning is the inference of latent and unknown parameters governing the physical system from observed data. This has typically been addressed through a variety of methods, including Physics-Informed Neural Networks (PINN) [17, 20] when there is a presumed description via differential equations, neural ODE systems [7] which embeds neural networks within ODE solvers, and Bayesian neural networks that could incorporate known mathematical models (e.g. B-PINN [25]).
@@
Most recently, Evidential Physics-Informed Neural Networks (E-PINN) was proposed in [22], where principles of Evidential Deep Learning (EDL) [^1] Evidential Deep Learning [2, 21] is a framework where a higher-order set of priors are placed over likelihood functions for the outputs, and model training is then used to infer uncertainty estimates from closed-form functions of the learned prior hyperparameters. was blended with those of PINN to yield uncertainty estimates for the target outputs assumed to follow a set of PDEs. E-PINN [22] was conjectured to be able to generate uncertainty estimates also for the unknown PDE parameters (apart from the target output) by averaging the PDE residual term over a hypothetical distribution. A limitation of the work of [22] is that it only applies to a subclass of PDEs where its parameters can be separated out as linear terms in a certain form (as expressed in equation 3 of [22]). It was also not clear how one should specify the hypothetical distribution for the learnable PDE parameters.
@@
In our work here, we present a significant refinement of the framework of [22], in particular, proposing a novel approach for inferring the unknown parameters that is much more aligned with Bayesian statistics. We formulate a data-driven, information-theoretic approach to specify a prior distribution for the parameters. Unlike ordinary PINN or the work in [22], the loss weight of the PDE residual is not fixed a priori but is learned adaptively, evolving throughout training. This refinement requires a two-phase model training algorithm where model is first trained purely on data in the first phase, allowing us to construct priors and appropriate initial conditions for the PDE residual and the learnable parameters, before invoking the PDE constraints in the second phase.
@@
As validation case studies, we applied our model to a 1D Poisson equation with Gaussian source and a 1+1D nonlinear Fisher-KPP equation, a prototype of the reaction-diffusion model. We compared the performance of our model against two popular uncertainty-aware frameworks that could also be synthesized with PINN. As reviewed in [8], a state-of-the-art proposal for uncertainty quantification in PINNs is the framework of Bayesian-PINN (B-PINN) of [25], where the PDE residual loss forms part of the posterior distribution. Another popular framework is that of Deep Ensemble ([11, 18]) where a set of similar models with different initial weights are taken to provide a distribution of predictions. The main performance metric we adopt here is the degree of calibration of the uncertainty quantification measured by the empirical coverage probability [12]. To illustrate its applicability in a real-world setting, we also apply E-PINN on published intravenous glucose tolerance test (IVGTT) data assuming the Bergman minimal model of glucose-insulin dynamics [3, 4] which has played a crucial role in modeling diabetes pathophysiology [5].
@@ 2.1 A posterior density from merging principles of Deep Evidential Regression and PINN
Now we take $P(\mathcal{D}|\mathcal{M}(\vec{w}))P(\mathcal{M}(\vec{w})|\vec{\Omega})$ as a joint likelihood function for $\vec{\Omega}$ which defines a posterior distribution for $\vec{\Omega}$ following eqn.(3). For the data loss term, we take it to be the term proposed in ‘Deep Evidential Regression’ [2] which maps a generic input variable to four output functions $\{\gamma,\alpha,\beta,\nu\}$. It is a t-distribution with density function
@@ 2.2 On uncertainty of $\vec{\Omega}$ and its prior distribution
where $N$ is the normalization constant and all integrals are implemented as numerical Riemann sums over the discretized domain $\mathcal{D}_{\vec{\Omega}}$. We would like the prior distribution of $\Omega$ to be characterized by the same mode and dispersion scales as the highest density region (HDR) of $f$ as discussed in [14]. At some confidence level, say 68%, this region is generally a complex subset of the domain $\mathcal{D}_{\Omega}$. Since our choice of prior distribution affects model training dynamics in the second phase, we adopt a simple Gaussian surrogate distribution for this region, with the means being the modes and the standard deviations being those of each marginal distribution.
@@ 2.3 Determination of $\pi(\sigma^{2}_{R};\alpha_{r},\beta_{r})$
$P(\mathcal{M}_{0}(\vec{w}^{0})|\vec{\Omega};\sigma^{2})=\frac{\text{exp}\left[-\frac{1}{2\sigma^{2}}\sum^{N_{D}}_{k=1}\mathcal{R}^{2}_{k}\left(\partial f,f,x_{k},\vec{\Omega}\right)\right]}{\int d\vec{\Omega}\,\,\,\text{exp}\left[-\frac{1}{2\sigma^{2}}\sum^{N_{D}}_{k=1}\mathcal{R}^{2}_{k}\left(\partial f,f,x_{k},\vec{\Omega}\right)\right]}$ (15)
@@ 2.4 A summary of model implementation
with each factor defined in eqns.(2), (6), (10) and (12). This second phase of training refines the purely data-fitted model such that it conforms to the presumed PDE description. Upon completion, we can obtain confidence bands on the model predictions, and infer $\vec{\Omega}$ with its uncertainty as defined by the mean and variance of the posterior distribution (3).
@@ 3 Validation case studies: 1D Poisson and 2D Fisher-KPP equations
We validated the E-PINN framework using two representative differential equations : (i)the 1D Poisson equation with Gaussian source (ii)the Fisher-KPP equation. These prototype equations and their variants have found wide applicability across scientific disciplines (see e.g. [6, 10, 13, 19].) For the main performance metric, we adopt the empirical coverage probability (ECP) to measure how well-calibrated the uncertainties are, relative to the nominal ones (NCP). In our case studies, we computed plots of ECP vs NCP and calculated the mean of the absolute discrepancy between these two quantities as a measure of the mean calibration error (MCE) (see e.g. [12, 15]). For assessing if the uncertainty in the inferred parameters $\vec{\Omega}$ is consistent with the model’s deviation from the PDE numerical solutions, we performed a Monte Carlo goodness-of-fit test (see e.g. [24]). This involves estimating an empirical p-value defined as the proportion of posterior samples whose deviation from the posterior mean exceeds that of the model curve.
@@
where $r,D$ are parameters to be inferred. They physically represent the growth and diffusion rates respectively when this equation is used to model reaction-diffusion systems that can be applied to wave propagation and population growth dynamics, etc. We construct our synthetic dataset based on an exact traveling wave solution first presented by Ablowitz and Zeppetella in [1].
@@ 3.3 Comparison against other methods of uncertainty quantification frameworks in PINN
Deep Ensemble method was implemented with 100 ordinary PINNs each formulated on the same 2-layer fully-connected perceptron trained with the ordinary mean-squared error loss function and the typical PINN residual loss function. The ensemble of models was defined by a random distribution in each model’s initial weights. Bayesian PINNs (B-PINN) were constructed following the main principles presented in [25]. Perceptron weights and PDE parameters were treated as latent variables, and inferred using Markov-Chain-Monte-Carlo sampling (implemented in NumPyro and Wax). Further we employed the No-U-Turn Sampler (NUTS), an adaptive variant of Hamiltonian Monte Carlo, which automatically adjusts the step size during warm-up to achieve efficient sampling and a desirable acceptance rate. For all models, training was performed on NVIDIA A100 GPU (40GB). The average times taken for the completion of training for the Poisson, Fisher-KPP case studies were: (i) (160 min, 56 min) for E-PINN (ii) (70 min, 100 min) for B-PINN with $10^{4}$ posterior samples with 30 percent burn-in (iii) (4500 min, 4000 min) for Deep Ensemble.
@@ 4 Application of E-PINN to Bergman minimal model of glucose-insulin regulatory system
In this Section, we present a brief case study of E-PINN modeling published clinical data [9, 16, 23] related to the intravenous glucose tolerance test (IVGTT) – a procedure that has been widely used in diabetes-related physiological research [5]. Within the literature related to mathematical models of glucose-insulin interactions, many works have been constructed based on a model due to Bergman et al. [3, 4] which is the following set of coupled first-order differential equations representing glucose-insulin dynamics.
@@
where $G$ represents glucose level, $I$ represents insulin concentration and $X$ a more abstract quantity representing glucose uptake activity of insulin-excitable tissue (e.g. liver, adipose, etc.). $G_{b},I_{b}$ are the basal concentrations of glucose and insulin representing baseline values when the effect of the bolus injection of glucose in IVGTT disappears. The parameters $p_{1},p_{2},p_{3}$ represent the glucose decay rate, spontaneous decay rate of remote insulin and the insulin-dependent increase in tissue glucose uptake ability respectively. These are the parameters we leverage E-PINN to infer from measurements of glucose and insulin evolutions in IVGTT samples. In particular, $p_{1}$ is known as an index measuring ‘glucose effectiveness’ (often denoted as $S_{G}$), while $S_{I}\equiv p_{3}/p_{2}$ is known as the insulin sensitivity index that measures the effectiveness at which glucose decreases upon release of insulin. In a IVGTT test, glucose and insulin samples are periodically collected over a period of a few hours. Taking the insulin to be an externally known function, the equations (23), (24) thus constitute a coupled set of first-order ODEs in the variables $G$ and $X$ which is often alluded to as the Bergman minimal model [3].
We apply E-PINN to the modeling of (23) and (24), using the datasets gathered and analyzed in [9, 16, 23]. In particular, we focused on modeling two datasets pertaining to the measured glucose and insulin levels corresponding to healthy and type 2 diabetic subjects. E-PINN enables probabilistic modeling of the datasets and directly yields the posterior distribution of the unknown parameters. Unlike the controlled case studies, we do not have empirical data points for the variable $X(t)$ and thus, our base neural network was designed to yield five outputs – four corresponding to $G(t)$ and its uncertainty, and another corresponding to $X(t)$ so as to avoid overfitting on $X(t)$. We define the PDE residual as the sum of the residual arising from (23) and (24) separately. Estimating $I_{b},G_{b}$ from the datasets, we are left with only $p_{1},p_{2},p_{3}$ as the unknown parameters to be inferred. Fig. 4 depicts our model’s prediction curves and the learned posterior distributions of two clinically useful indices: the glucose effectiveness index $S_{G}$ and insulin sensitivity index $S_{I}$, the latter appearing to be the more discriminatory one, being about 3 orders of magnitude larger for healthy individuals.
@@ 5 Conclusion
We have constructed a class of PINN that is equipped with an uncertainty quantification framework for its target and PDE parameter inferences. It leverages the marginal distribution-based loss function of EDL for estimating uncertainty of outputs and infers unknown parameters of the PDE through learning a posterior distribution. Two key features distinguish our model from the usual PINN: (i)the loss weight of the PDE residual loss term evolves dynamically throughout model training, (ii)a prior distribution based on a purely data-fitted model regularizes the learning of the unknown PDE parameters. Validating our model on two controlled case studies involving the 1D Poisson equation with a Gaussian source and the 2D Fisher-KPP equation, we found that E-PINN generated empirical coverage probabilities that were calibrated significantly better than the B-PINN and Deep Ensemble methods. It inferred the unknown parameters accurately within the bounds of confidence intervals consistent with the Monte-Carlo goodness-of-fit test. In the Fisher-KPP case, E-PINN’s uncertainty distribution correlated well with a temporally varying noise, unlike the other two frameworks. As an illustrative real-world application, we demonstrated how E-PINN can be applied to model clinical IVGTT datasets presumed to follow Bergman minimal model of glucose-insulin dynamics, with results indicating that the insulin sensitivity index $S_{I}$ discriminates effectively between healthy and diabetic patients. We look forward to more extensive validation and application of E-PINN to solving inverse problems across scientific disciplines.

### v3

@@ 1 Introduction
A central problem in scientific machine learning is the inference of latent and unknown parameters governing the physical system from observed data. This has typically been addressed through a variety of methods, including Physics-Informed Neural Networks (PINN) [Lagaris_1998, pinn] when there is a presumed description via differential equations, neural ODE systems [neural] which embeds neural networks within ODE solvers, and Bayesian neural networks that could incorporate known mathematical models (e.g. B-PINN [bpinn]).
@@
Most recently, Evidential Physics-Informed Neural Networks (E-PINN) was proposed in [tan3], where principles of Evidential Deep Learning (EDL) [^1] Evidential Deep Learning [amini, sensoy] is a framework where a higher-order set of priors are placed over likelihood functions for the outputs, and model training is then used to infer uncertainty estimates from closed-form functions of the learned prior hyperparameters. was blended with those of PINN to yield uncertainty estimates for the target outputs assumed to follow a set of PDEs. E-PINN [tan3] was conjectured to be able to generate uncertainty estimates also for the unknown PDE parameters (apart from the target output) by averaging the PDE residual term over a hypothetical distribution. A limitation of the work of [tan3] is that it only applies to a subclass of PDEs where its parameters can be separated out as linear terms in a certain form (as expressed in equation 3 of [tan3]). It was also not clear how one should specify the hypothetical distribution for the learnable PDE parameters.
@@
In our work here, we present a significant refinement of the framework of [tan3], in particular, proposing a novel approach for inferring the unknown parameters that is much more aligned with Bayesian statistics. We formulate a data-driven, information-theoretic approach to specify a prior distribution for the parameters. Unlike ordinary PINN or the work in [tan3], the loss weight of the PDE residual is not fixed a priori but is learned adaptively, evolving throughout training. This refinement requires a two-phase model training algorithm where model is first trained purely on data in the first phase, allowing us to construct priors and appropriate initial conditions for the PDE residual and the learnable parameters, before invoking the PDE constraints in the second phase.
@@
As validation case studies, we applied our model to a 1D Poisson equation with Gaussian source and a 1+1D nonlinear Fisher-KPP equation, a prototype of the reaction-diffusion model. We compared the performance of our model against two popular uncertainty-aware frameworks that could also be synthesized with PINN. As reviewed in [cuomo2022], a state-of-the-art proposal for uncertainty quantification in PINNs is the framework of Bayesian-PINN (B-PINN) of [bpinn], where the PDE residual loss forms part of the posterior distribution. Another popular framework is that of Deep Ensemble ([Ganaie, lak]) where a set of similar models with different initial weights are taken to provide a distribution of predictions. The main performance metric we adopt here is the degree of calibration of the uncertainty quantification measured by the empirical coverage probability [ECP]. To illustrate its applicability in a real-world setting, we also apply E-PINN on published intravenous glucose tolerance test (IVGTT) data assuming the Bergman minimal model of glucose-insulin dynamics [bergman2, bergman1] which has played a crucial role in modeling diabetes pathophysiology [bergman3].
@@ 2.1 A posterior density from merging principles of Deep Evidential Regression and PINN
Now we take $P(\mathcal{D}|\mathcal{M}(\vec{w}))P(\mathcal{M}(\vec{w})|\vec{\Omega})$ as a joint likelihood function for $\vec{\Omega}$ which defines a posterior distribution for $\vec{\Omega}$ following eqn.(3). For the data loss term, we take it to be the term proposed in ‘Deep Evidential Regression’ [amini] which maps a generic input variable to four output functions $\{\gamma,\alpha,\beta,\nu\}$. It is a t-distribution with density function
@@ 2.2 On uncertainty of $\vec{\Omega}$ and its prior distribution
where $N$ is the normalization constant and all integrals are implemented as numerical Riemann sums over the discretized domain $\mathcal{D}_{\vec{\Omega}}$. We would like the prior distribution of $\Omega$ to be characterized by the same mode and dispersion scales as the highest density region (HDR) of $f$ as discussed in [Hyndman]. At some confidence level, say 68%, this region is generally a complex subset of the domain $\mathcal{D}_{\Omega}$. Since our choice of prior distribution affects model training dynamics in the second phase, we adopt a simple Gaussian surrogate distribution for this region, with the means being the modes and the standard deviations being those of each marginal distribution.
@@ 2.3 Determination of $\pi(\sigma^{2}_{R};\alpha_{r},\beta_{r})$
$P(\mathcal{M}_{0}(\vec{w}^{0})|\vec{\Omega};\sigma^{2})=\frac{\text{exp}\left[-\frac{1}{2\sigma^{2}}\sum^{N_{D}}_{k=1}\mathcal{R}^{2}_{k}\left(\partial f,f,x_{k},\vec{\Omega}\right)\right]}{\int d\vec{\Omega}\,\,\,\text{exp}\left[-\frac{1}{2\sigma^{2}}\sum^{N_{D}}_{k=1}\mathcal{R}^{2}_{k}\left(\partial f,f,x_{k},\vec{\Omega}\right)\right]}.$ (15)
@@ 2.4 A summary of model implementation
with each factor defined in eqns.(2), (6), (10) and (12). This second phase of training refines the purely data-fitted model such that it conforms to the presumed PDE description. Upon completion, we can obtain confidence bands on the model predictions, and infer $\vec{\Omega}$ with uncertainty as defined by the mean and variance of the posterior (3). Our methodology is broadly applicable to systems of differential equations, irrespective of the choice of differential operators or dimensionality, as reflected in the nature of eqns. (2), (6), (10) and (12).
@@ 3 Validation case studies: 1D Poisson and 2D Fisher-KPP equations
We validated the E-PINN framework using two representative differential equations : (i)the 1D Poisson equation with Gaussian source (ii)the Fisher-KPP equation. These prototype equations and their variants have found wide applicability across scientific disciplines (see e.g. [Poisson1, Fisher1, Fisher2, Poisson2].) For the main performance metric, we adopt the empirical coverage probability (ECP) to measure how well-calibrated the uncertainties are, relative to the nominal ones (NCP). In our case studies, we computed plots of ECP vs NCP and calculated the mean of the absolute discrepancy between these two quantities as a measure of the mean calibration error (MCE) (see e.g. [ECP, jungo]). For assessing if the uncertainty in the inferred parameters $\vec{\Omega}$ is consistent with the model’s deviation from the PDE numerical solutions, we performed a Monte Carlo goodness-of-fit test (see e.g. [weiss]). This involves estimating an empirical p-value defined as the proportion of posterior samples whose deviation from the posterior mean exceeds that of the model curve.
@@
where $r,D$ are parameters to be inferred. They physically represent the growth and diffusion rates respectively when this equation is used to model reaction-diffusion systems that can be applied to wave propagation and population growth dynamics, etc. We construct our synthetic dataset based on an exact traveling wave solution first presented by Ablowitz and Zeppetella in [ablo].
@@ 3.3 Comparison against other methods of uncertainty quantification frameworks in PINN
Deep Ensemble method was implemented with 100 ordinary PINNs each formulated on the same 2-layer fully-connected perceptron trained with the ordinary mean-squared error loss function and the typical PINN residual loss function. The ensemble of models was defined by a random distribution in each model’s initial weights. Bayesian PINNs (B-PINN) were constructed following the main principles presented in [bpinn]. Perceptron weights and PDE parameters were treated as latent variables, and inferred using Markov-Chain-Monte-Carlo sampling (implemented in NumPyro and Wax). Further we employed the No-U-Turn Sampler (NUTS), an adaptive variant of Hamiltonian Monte Carlo, which automatically adjusts the step size during warm-up to achieve efficient sampling and a desirable acceptance rate. For all models, training was performed on NVIDIA A100 GPU (40GB). In the aspect of computational cost, the average times taken for the completion of training for the Poisson, Fisher-KPP case studies were: (i) (160 min, 56 min) for E-PINN (ii) (70 min, 100 min) for B-PINN with $10^{4}$ posterior samples with 30 percent burn-in (iii) (4500 min, 4000 min) for Deep Ensemble.
@@ 4 Application of E-PINN to Bergman minimal model of glucose-insulin regulatory system
In this Section, we present a brief case study of E-PINN modeling published clinical data [Gaetano, Kartono, Ward] related to the intravenous glucose tolerance test (IVGTT) – a procedure that has been widely used in diabetes-related physiological research [bergman3]. Within the literature related to mathematical models of glucose-insulin interactions, many works have been constructed based on a model due to Bergman et al. [bergman2, bergman1] which is the following set of coupled first-order differential equations representing glucose-insulin dynamics.
@@
where $G$ represents glucose level, $I$ represents insulin concentration and $X$ a more abstract quantity representing glucose uptake activity of insulin-excitable tissue (e.g. liver, adipose, etc.). $G_{b},I_{b}$ are the basal concentrations of glucose and insulin representing baseline values when the effect of the bolus injection of glucose in IVGTT disappears. The parameters $p_{1},p_{2},p_{3}$ represent the glucose decay rate, spontaneous decay rate of remote insulin and the insulin-dependent increase in tissue glucose uptake ability respectively. These are the parameters we leverage E-PINN to infer from measurements of glucose and insulin evolutions in IVGTT samples. In particular, $p_{1}$ is known as an index measuring ‘glucose effectiveness’ (often denoted as $S_{G}$), while $S_{I}\equiv p_{3}/p_{2}$ is known as the insulin sensitivity index that measures the effectiveness at which glucose decreases upon release of insulin. In a IVGTT test, glucose and insulin samples are periodically collected over a period of a few hours. Taking the insulin to be an externally known function, the equations (23), (24) thus constitute a coupled set of first-order ODEs in the variables $G$ and $X$ which is often alluded to as the Bergman minimal model [bergman2].
We apply E-PINN to the modeling of (23) and (24), using the datasets gathered and analyzed in [Gaetano, Kartono, Ward]. In particular, we focused on modeling two datasets pertaining to the measured glucose and insulin levels corresponding to healthy and type 2 diabetic subjects. E-PINN enables probabilistic modeling of the datasets and directly yields the posterior distribution of the unknown parameters. Unlike the controlled case studies, we do not have empirical data points for the variable $X(t)$ and thus, our base neural network was designed to yield five outputs – four corresponding to $G(t)$ and its uncertainty, and another corresponding to $X(t)$ so as to avoid overfitting on $X(t)$. We define the PDE residual as the sum of the residual arising from (23) and (24) separately. Estimating $I_{b},G_{b}$ from the datasets, we are left with only $p_{1},p_{2},p_{3}$ as the unknown parameters to be inferred. Fig. 4 depicts our model’s prediction curves and the learned posterior distributions of two clinically useful indices: the glucose effectiveness index $S_{G}$ and insulin sensitivity index $S_{I}$, the latter appearing to be the more discriminatory one, being about 3 orders of magnitude larger for healthy individuals.
@@ 5 Conclusion
We have constructed a class of PINN that is equipped with an uncertainty quantification framework for its target and PDE parameter inferences. It leverages the loss function of EDL for estimating uncertainty of outputs and infers unknown PDE parameters through learning a posterior distribution. Two key features distinguish our model from the usual PINN: (i)the loss weight of the PDE residual term evolves throughout training, (ii)a prior distribution based on a purely data-fitted model regularizes learning of the unknown parameters. Our methodology is broadly applicable to systems of differential equations, irrespective of the choice of differential operators or dimensionality. Validating our model on two controlled case studies involving the 1D Poisson equation with a Gaussian source and the 2D Fisher-KPP equation, we found that E-PINN generated empirical coverage probabilities that were calibrated significantly better than the B-PINN and Deep Ensemble methods. It inferred unknown parameters accurately within the bounds of confidence intervals consistent with the Monte-Carlo goodness-of-fit test. In the Fisher-KPP case, E-PINN’s uncertainty distribution correlated well with a temporally varying noise, unlike the other two frameworks. As an illustrative real-world application, we demonstrated how E-PINN can be applied to model clinical IVGTT datasets presumed to follow Bergman minimal model of glucose-insulin dynamics, with results indicating that the insulin sensitivity index $S_{I}$ discriminates effectively between healthy and diabetic patients. We look forward to extensive applications of E-PINN to solving inverse problems across scientific disciplines.